# Latent Diffusion Model without Variational Autoencoder

### v3

@@ Abstract
Recent progress in diffusion-based visual generation has largely relied on latent diffusion models with Variational Autoencoders (VAEs). While effective for high-fidelity synthesis, this VAE+Diffusion paradigm suffers from limited training efficiency, slow inference, and poor transferability to broader vision tasks. These issues stem from a key limitation of VAE latent spaces: the lack of clear semantic separation and strong discriminative structure. Our analysis confirms that these properties are not only crucial for perception and understanding tasks, but also equally essential for the stable and efficient training of latent diffusion models. Motivated by this insight, we introduce SVG—a novel latent diffusion model without variational autoencoders, which unleashes Self-supervised representations for Visual Generation. SVG constructs a feature space with clear semantic discriminability by leveraging frozen DINO features, while a lightweight residual branch captures fine-grained details for high-fidelity reconstruction. Diffusion models are trained directly on this semantically structured latent space to facilitate more efficient learning. As a result, SVG enables accelerated diffusion training, supports few-step sampling, and improves generative quality. Experimental results further show that SVG preserves the semantic and discriminative capabilities of the underlying self-supervised representations, providing a principled pathway toward task-general, high-quality visual representations.
@@
Figure 1: Core contribution of SVG. (a-d) Comparisons of the overall methodology between VAE-based latent diffusion models and SVG. (e-f) Comparisons of training and inference efficiency.
@@ 1 Introduction
Despite their success, the VAE+Diffusion paradigm exhibits several critical limitations. First, both training and inference are computationally expensive: for instance, training an ImageNet $256\times 256$ generation model with the standard DiT implementation (Peebles & Xie, 2022) requires 7M steps, and inference typically demands more than 25 sampling steps to achieve satisfactory results. As depicted in parts (a)-(c) of Figure 1, although some recent methods (Yu et al., 2025; Leng et al., 2025; Yao et al., 2025) attempt to accelerate diffusion training by aligning with external feature spaces of vision foundation models (VFM) (Oquab et al., 2023; He et al., 2021; Chen et al., 2020d; Radford et al., 2021; Zhai et al., 2023) or imposing regularization constraints on the VAE latent space (Wang & He, 2025; Stoica et al., 2025), these approaches provide only ad hoc fixes, as they do not fundamentally alter the VAE training objective or the resulting latent space structure, which inherently lacks semantic separability. Importantly, VAE latent representations are generally not employed in modern multi-modal large models, and their restricted perceptual capabilities (Yin et al., 2024; Jin et al., 2024) highlight a fundamental limitation. This discrepancy implies that VAE latents are unlikely to serve effectively as unified visual representations.
In this paper, we argue that a discriminative semantic structure in the latent space can substantially facilitate the training of diffusion models. By leveraging the powerful self-supervised features from DINOv3, we demonstrate that it is possible to construct a feature space that enables efficient diffusion training in a simple yet effective way, while fully retaining DINOv3’s strengths beyond generation.
@@ 2 Related works
Visual generation. Generative models aim to learn the underlying probability distribution of data and to generate novel samples that are both realistic and diverse. Generative adversarial networks (GANs) (Goodfellow et al., 2014; Radford et al., 2016; Arjovsky et al., 2017; Gulrajani et al., 2017; Karras et al., 2019; Zhu et al., 2017; Karras et al., 2018; Sauer et al., 2022) generate realistic images via adversarial training but often suffer from mode collapse, instability, and poor interpretability. Another family of approaches follows an autoregressive paradigm, where an image is represented as a sequence of pixels, patches, or latent tokens. The joint distribution is factorized into conditional probabilities and modeled sequentially, as in (Salimans et al., 2017; Vaswani et al., 2018; Chen et al., 2020a). Extensions based on masked autoregression (He et al., 2021; Chang et al., 2022; Li et al., 2024) predict missing tokens given visible context, analogous to masked language models in NLP. This formulation enables direct transfer of transformer-based sequence modeling techniques to large-scale image generation. More recently, diffusion models (Ho et al., 2020; Nichol & Dhariwal, 2021; Song et al., 2021a; b) have emerged as a powerful alternative, generating images by iteratively denoising Gaussian noise. They achieve state-of-the-art fidelity and diversity, with improved training stability and mode coverage. An improved version, the latent diffusion model (LDM) (Rombach et al., 2021; Peebles & Xie, 2022; Ma et al., 2024; Liu et al., 2022), integrates a VAE (Kingma & Welling, 2022) with the diffusion process to operate in a lower-dimensional latent space, reducing computational cost while maintaining generation quality. Nevertheless, these models still require multiple inference steps, limiting generation speed. Beyond these core paradigms, recent research (Yu et al., 2025; Leng et al., 2025; Yao et al., 2025; Li et al., 2023b) has also explored leveraging external visual representations—such as pre-trained feature extractors or vision-language models—to enhance the efficiency, controllability, and overall quality of LDMs, highlighting the increasing trend of integrating generative modeling with strong discriminative representations. However, we argue that such feature alignment still cannot overcome the inherent limitations of the VAE+Diffusion paradigm, resulting in constrained training and inference efficiency, and hindering the unified feature modeling of visual generation, perception, and understanding.
@@
Visual representation learning. Recent advances in visual representation learning can be broadly categorized into discriminative, generative, and multimodal paradigms. Disciminative methods, such as self-supervised learning (SSL) approaches including DINO (Zhang et al., 2022; Oquab et al., 2023; Siméoni et al., 2025), SimCLR (Chen et al., 2020b; c), MoCo (He et al., 2019; Chen et al., 2020d), and BYOL (Grill et al., 2020), learn informative features without explicitly modeling the data distribution, producing highly separable representations that excel in classification, retrieval, and dense prediction, but often discard fine grained generative information, limiting their utility for synthesis or reconstruction. Generative approaches, including VAE (Kingma & Welling, 2022), masked autoencoders (MAE) (He et al., 2021), masked image modeling (MIM) (Xie et al., 2022b), and diffusion models (Ho et al., 2020; Song et al., 2021b), capture rich contextual and perceptual information by reconstructing inputs or modeling the underlying data distribution, providing embeddings beneficial for downstream tasks; however, they are computationally intensive and may produce representations that are less discriminative than contrastive methods. Multimodal methods, exemplified by CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023; Tschannen et al., 2025), Florence (Xiao et al., 2023), and BLIP (Li et al., 2022; 2023a), align images and text in a shared latent space, enabling zero-shot learning, cross-modal retrieval, and enhanced semantic understanding, but rely on large-scale paired data and can underperform when one modality dominates or data quality is uneven. Despite these advances, existing visual representations often struggle to provide a unified solution for various visual tasks. Here, we for the first time demonstrate that features learned via self-supervised methods can be directly repurposed for generative modeling, enabling the construction of a unified feature space that effectively supports diverse core vision tasks.
@@ 3.1 Preliminaries
${\mathcal{L}}_{\mathrm{FM}}=\mathbb{E}_{{\mathbf{x}}_{0}\sim p_{0}({\mathbf{x}}),{\mathbf{\epsilon}}\sim p_{1}({\mathbf{x}})}[\lambda(t)\|{\mathbf{v}}_{\theta}({\mathbf{x}}_{t},t)-{\mathbf{v}}_{t}].$ (5)
@@ 3.2 Rethinking Latent Diffusion Models
Figure 4: Visualization of feature space. (a) Feature visualization with t-SNE for 100 ImageNet classes (100 random samples per class, top row) and 20 classes (100 random samples per class, bottom row). Features are extracted using DINOv3 (Siméoni et al., 2025), VA-VAE (Yao et al., 2025), SD-VAE (Rombach et al., 2021), and MAR-VAE (Li et al., 2024), with each class shown in a distinct color. (b) Each subfigure shows the source distribution (black dots) and the target distribution, where the samples are divided into two semantic categories (green and blue dots). The arrows indicate the directions of the mean velocity field at each point.
3.3 Feature Visual Generation
@@
SVG diffusion. Unlike prior approaches that construct diffusion models on low-dimensional VAE latent spaces (Rombach et al., 2021), SVG Diffusion treats the high-dimensional SVG feature space as the target distribution, trained using the flow matching objective defined in Equation 5. Specifically, for $256\times 256$ images, the DINOv3-ViT-S/16+ encoder produces a $16\times 16\times 384$ feature map, compared with the $16\times 16\times 4$ VAE latent in DiT (Peebles & Xie, 2022). While training diffusion models in such high-dimensional spaces is generally challenging and prone to unstable convergence (Xie et al., 2024), the well-dispersed semantic structure of SVG features makes training stable and efficient. Consequently, SVG Diffusion converges faster and achieves superior generative quality compared with VAE-based diffusion. Moreover, since hidden states in diffusion models typically have channel sizes larger than 384 (Peebles & Xie, 2022), SVG Diffusion does not incur inference inefficiency. As shown in Section 4.3, its strong semantic continuity also enables few-step sampling, yielding superior inference efficiency.
SVG training pipeline. The training is conducted in two stages. In the first stage, we optimize only the Residual Encoder and the SVG Decoder with the reconstruction loss defined in (Rombach et al., 2021). However, naively training in this way causes the decoder to over-rely on the Residual Encoder, and the mismatch in numerical ranges between the DINO and residual outputs can compromise the semantic discriminability inherited from DINO. To address this, we align the Residual Encoder outputs with the DINO feature distribution, ensuring that the added residual dimensions do not distort the original semantic space. In the second stage, SVG Diffusion is trained under the settings of SiT (Ma et al., 2024), with QK-Norm (Henry et al., 2020) applied and the per-channel SVG feature space normalized to stabilize training.
@@ 4.3 Analysis
Preserving the original capabilities of DINO features. The previous experiments have verified the superiority of the SVG space in visual generation. To test whether it also preserves visual perception and understanding capability, we evaluate the SVG encoder against the DINO encoder on downstream tasks where DINO is known to excel. For each task, we adopt a simple strategy: a lightweight MLP- or linear-layer decoder is appended to the encoder to map features into predictions, and only the decoder is trained (details in Appendix B). As reported in Table 4, the SVG feature maintains the strong generalization ability of DINO, achieving comparable or even slightly superior results on ImageNet-1K (Deng et al., 2009; Russakovsky et al., 2015) classification, ADE20K (Zhou et al., 2019) semantic segmentation, and NYUv2 (Nathan Silberman & Fergus, 2012) depth estimation. Combined with its previously demonstrated strength in generative tasks, this dual advantage establishes the feature space produced by  SVG Encoder as a unified representation space for diverse vision tasks.
@@ Appendix A More implementation details
Classifier-free guidance. In the main text, we report results both with and without classifier-free guidance. To reduce the uncertainty in the initial velocity prediction, we adopt zero-init (Fan et al., 2025), which skips the first step. We report the FID50K with cfg 1.55 for SVG in our paper.
@@ Appendix B Finetuning Details on Downstream Tasks
We evaluate the SVG encoder against the DINOv3 encoder on three representative downstream tasks: ImageNet-1K (Deng et al., 2009; Russakovsky et al., 2015) classification, ADE20K (Zhou et al., 2019) semantic segmentation, and NYUv2 (Nathan Silberman & Fergus, 2012) depth estimation. For all tasks, the encoder remains frozen, and only lightweight decoders are trained.
Image classification. This task requires assigning each image to a single class. We report Top-1 and Top-5 accuracies. A linear classifier is placed on top of the encoder output to map features to class scores. Input images are randomly resized and cropped to $256\times 256$. The classifier is trained for 30 epochs with the AdamW optimizer (Loshchilov & Hutter, 2017), using a global batch size of 3072, an initial learning rate of 5e45e-4, and weight decay of 1e21e-2.
Semantic segmentation. The goal of semantic segmentation is to produce dense per-pixel predictions. We use mean Intersection-over-Union (mIoU) and mean Accuracy (mAcc) as evaluation metrics. The decoder adopts an FPNHead implementation in mmsegmentation, applied to the single-scale encoder feature. The training strategy generally follows (Zhao et al., 2023). Images are randomly resized and cropped to $512\times 512$ before being fed to the network. Optimization is performed with AdamW (Loshchilov & Hutter, 2017) at a learning rate of 8e58e-5, weight decay of 1e31e-3, and 1,500 warm-up steps. A polynomial scheduler with power 0.9 and a minimum learning rate of 1e61e-6 is used. Training runs for 8,000 iterations. During inference, we adopt sliding-window evaluation with $512\times 512$ crops and a stride of $341\times 341$.
Depth estimation. Depth estimation aims to regress pixel-wise depth values for input images. We report Absolute Relative Error (A.Rel) and RMSE as evaluation metrics. During training, images are randomly cropped to $480\times 480$. The model is optimized for 25 epochs using the AdamW (Loshchilov & Hutter, 2017) optimizer with a batch size of 24 and a learning rate of 5e45e-4. The decoder head and other hyperparameters follow the setup in (Xie et al., 2022a). At test time, we use both horizontal flipping and sliding-window inference.

### v4

@@ Abstract
Recent progress in diffusion-based visual generation has largely relied on latent diffusion models with Variational Autoencoders (VAEs). While effective for high-fidelity synthesis, this VAE+Diffusion paradigm still suffers from limited training and inference efficiency, along with poor transferability to broader vision tasks. These issues stem from a key limitation of VAE latent spaces: the lack of clear semantic separation and strong discriminative structure. Our analysis confirms that these properties are not only crucial for perception and understanding tasks, but also equally essential for the stable and efficient training of latent diffusion models. Motivated by this insight, we introduce SVG—a novel latent diffusion model without variational autoencoders, which unleashes Self-supervised representations for Visual Generation. SVG constructs a feature space with clear semantic discriminability by leveraging frozen DINO features, while a lightweight residual branch captures fine-grained details for high-fidelity reconstruction. Diffusion models are trained directly on this semantically structured latent space to facilitate more efficient learning. As a result, SVG enables accelerated diffusion training, supports few-step sampling, and improves generative quality. Experimental results further show that SVG preserves the semantic and discriminative capabilities of the underlying self-supervised representations, providing a principled pathway toward task-general, high-quality visual representations.
@@
Figure 1: Core contribution of SVG. (a) Vanilla VAE-based LDM: the diffusion model is trained on the pretrained VAE latent space. (b) Diff. Model Feature Alignment: intermediate features of the diffusion model are aligned to Visual Foundation Model (VFM) features. (c) VAE and Diff. Model Feature Alignment: both VAE latent features and diffusion model intermediate features are aligned to VFM features. (d) Our method: the diffusion model is trained directly in the SVG space derived from self-supervised representations (DINOv3). (e-f) Comparisons of inference and training efficiency.
@@ 1 Introduction
Despite their success, the VAE+Diffusion paradigm exhibits several critical limitations. First, both training and inference are computationally expensive: for instance, training an ImageNet $256\times 256$ generation model with the standard DiT implementation (Peebles and Xie, 2022) requires 7M steps, and inference typically demands more than 25 sampling steps to achieve satisfactory results. As depicted in parts (a)-(c) of Figure 1, although some recent methods (Yu et al., 2025; Leng et al., 2025; Yao et al., 2025) attempt to accelerate diffusion training by aligning with external feature spaces of vision foundation models (VFM) (Oquab et al., 2023; He et al., 2021; Chen et al., 2020d; Radford et al., 2021; Zhai et al., 2023) or imposing regularization constraints on the VAE latent space (Wang and He, 2025; Stoica et al., 2025), these approaches provide only ad hoc fixes, as they do not fundamentally alter the semantic structure of the latent space, which remain weakly discriminable as clearly shown in Figure 4(a). Importantly, VAE latent representations are generally not employed in modern multi-modal large models, and their restricted perceptual capabilities (Yin et al., 2024; Jin et al., 2024) highlight a fundamental limitation. This discrepancy implies that VAE latents are unlikely to serve effectively as unified visual representations.
In this paper, we argue that a discriminative semantic structure in the latent space can substantially facilitate the training of diffusion models. By leveraging the powerful self-supervised features from DINOv3 (Siméoni et al., 2025), we demonstrate that it is possible to construct a feature space that enables efficient diffusion training in a simple yet effective way, while fully retaining DINOv3’s strengths beyond generation.
@@ 2 Related works
Visual generation. Generative models aim to learn the underlying probability distribution of data and to generate novel samples that are both realistic and diverse. Generative adversarial networks (GANs) (Goodfellow et al., 2014; Radford et al., 2016; Arjovsky et al., 2017; Gulrajani et al., 2017; Karras et al., 2019; Zhu et al., 2017; Karras et al., 2018; Sauer et al., 2022) generate realistic images via adversarial training but often suffer from mode collapse, instability, and poor interpretability. Another family of approaches follows an autoregressive paradigm, where an image is represented as a sequence of pixels, patches, or latent tokens. The joint distribution is factorized into conditional probabilities and modeled sequentially, as in (Salimans et al., 2017; Vaswani et al., 2018; Chen et al., 2020a). Extensions based on masked autoregression (He et al., 2021; Chang et al., 2022; Li et al., 2024) predict missing tokens given visible context, analogous to masked language models in NLP. This formulation enables direct transfer of transformer-based sequence modeling techniques to large-scale image generation. More recently, diffusion models (Ho et al., 2020; Nichol and Dhariwal, 2021; Song et al., 2021a; b) have emerged as a powerful alternative, generating images by iteratively denoising Gaussian noise. They achieve state-of-the-art fidelity and diversity, with improved training stability and mode coverage. An improved version, the latent diffusion model (LDM) (Rombach et al., 2021; Peebles and Xie, 2022; Ma et al., 2024; Liu et al., 2022), integrates a VAE (Kingma and Welling, 2022) with the diffusion process to operate in a lower-dimensional latent space, reducing computational cost while maintaining generation quality. Nevertheless, these models still require multiple inference steps, limiting generation speed. A line of research aims to improve the structural properties of the VAE latent space (Burgess et al., 2018; Xu and Durrett, 2018; Kouzelis et al., 2025a; Skorokhodov et al., 2025) or to incorporate diffusion processes into VAE reconstruction (Preechakul et al., 2022; Pandey et al., 2021; Vahdat et al., 2021). More recently, efforts have focused on aligning the latent features of diffusion models with external semantic representations (Yu et al., 2025; Leng et al., 2025; Yao et al., 2025; Li et al., 2023b), while other works explore joint generation approaches that integrate discriminative components into the generative process (Wu et al., 2025; Kouzelis et al., 2025b). Despite these advances, we observe that even with structural improvements to the VAE latent space or alignment with external semantic features, the semantic discriminability of VAE representations remains inherently limited. Consequently, the VAE+Diffusion paradigm is constrained in both training and inference efficiency, and it fails to establish a unified feature space capable of supporting visual generation, perception, and understanding.
@@
Visual representation learning. Recent advances in visual representation learning can be broadly categorized into discriminative, generative, and multimodal paradigms. Disciminative methods, such as self-supervised learning (SSL) approaches including DINO (Zhang et al., 2022; Oquab et al., 2023; Siméoni et al., 2025), SimCLR (Chen et al., 2020b; c), MoCo (He et al., 2019; Chen et al., 2020d), and BYOL (Grill et al., 2020), learn informative features without explicitly modeling the data distribution, producing highly separable representations that excel in classification, retrieval, and dense prediction, but often discard fine grained generative information, limiting their utility for synthesis or reconstruction. Generative approaches, including VAE (Kingma and Welling, 2022), masked autoencoders (MAE) (He et al., 2021), masked image modeling (MIM) (Xie et al., 2022b), and diffusion models (Ho et al., 2020; Song et al., 2021b), capture rich contextual and perceptual information by reconstructing inputs or modeling the underlying data distribution, providing embeddings beneficial for downstream tasks; however, they are computationally intensive and may produce representations that are less discriminative than contrastive methods. Multimodal methods, exemplified by CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023; Tschannen et al., 2025), Florence (Xiao et al., 2023), and BLIP (Li et al., 2022; 2023a), align images and text in a shared latent space, enabling zero-shot learning, cross-modal retrieval, and enhanced semantic understanding, but rely on large-scale paired data and can underperform when one modality dominates or data quality is uneven. Despite these advances, existing visual representations often struggle to provide a unified solution for various visual tasks. Here, we for the first time demonstrate that features learned via self-supervised methods can be directly repurposed for generative modeling, enabling the construction of a unified feature space that effectively supports diverse core vision tasks.
@@ 3.1 Preliminaries
${\mathcal{L}}_{\mathrm{FM}}=\mathbb{E}_{{\mathbf{x}}_{0}\sim p_{0}({\mathbf{x}}),{\mathbf{\epsilon}}\sim p_{1}({\mathbf{x}})}[\lambda(t)\|{\mathbf{v}}_{\theta}({\mathbf{x}}_{t},t)-{\mathbf{v}}_{t}\|].$ (5)
@@ 3.2 Rethinking Latent Diffusion Models
Figure 4: Visualization of feature space. (a) Feature visualization with t-SNE for 100 ImageNet classes (100 random samples per class, top row) and 20 classes (100 random samples per class, bottom row). Features are extracted using DINOv3 (Siméoni et al., 2025), VA-VAE (Yao et al., 2025), SD-VAE (Rombach et al., 2021), and MAR-VAE (Li et al., 2024), with each class shown in a distinct color, and model names annotated with their linear-probe Top-1 accuracy on ImageNet-1K (Deng et al., 2009). (b) Each subfigure shows the source distribution (black dots) and the target distribution, where the samples are divided into two semantic categories (green and blue dots). The arrows indicate the directions of the mean velocity field at each point.
3.3 Visual Feature Generation
@@
SVG diffusion. Unlike prior approaches that construct diffusion models on low-dimensional VAE latent spaces (Rombach et al., 2021), SVG Diffusion treats the high-dimensional SVG feature space as the target distribution, trained using the flow matching objective defined in Equation 5. Specifically, for $256\times 256$ images, the DINOv3-ViT-S/16+ encoder produces a $16\times 16\times 384$ feature map, compared with the $16\times 16\times 4$ VAE latent in DiT (Peebles and Xie, 2022). While training diffusion models in such high-dimensional spaces is generally challenging and prone to unstable convergence (Xie et al., 2024), the well-dispersed semantic structure of SVG features makes training stable and efficient. Consequently, SVG Diffusion converges faster and achieves superior generative quality compared with VAE-based diffusion. Moreover, since hidden states in diffusion models typically have channel sizes larger than 384 (Peebles and Xie, 2022), SVG Diffusion does not incur inference inefficiency. As shown in Section 4.3, its strong semantic continuity also enables few-step sampling, yielding superior inference efficiency.
SVG training pipeline. The training is conducted in two stages. In the first stage, we optimize only the Residual Encoder and the SVG Decoder with the reconstruction loss defined in (Rombach et al., 2021). However, naively training in this way causes the decoder to over-rely on the Residual Encoder, and the mismatch in numerical ranges between the DINO and residual outputs can compromise the semantic discriminability inherited from DINO. To address this, we align the Residual Encoder outputs with the DINO feature distribution, ensuring that the added residual dimensions do not distort the original semantic space. In specific, for each training batch, let $F_{D}$ denote the DINOv3 features and $F_{R}$ the residual features. The alignment is performed by normalizing $F_{R}$ to match the batch statistics of $F_{D}$:
$\hat{F}_{R}=\frac{F_{R}-\mu(F_{R})}{\sigma(F_{R})}\cdot\sigma(F_{D})+\mu(F_{D})$ (6)
where $\mu(\cdot)$ and $\sigma(\cdot)$ compute the mean and standard deviation along the feature dimension.
In the second stage, SVG Diffusion is trained under the settings of SiT (Ma et al., 2024), with QK-Norm (Henry et al., 2020) applied and the per-channel SVG feature space normalized to stabilize training.
@@ 4.3 Analysis
Preserving the original capabilities of DINO features. The previous experiments have verified the superiority of the SVG space in visual generation. To test whether it also preserves visual perception and understanding capability, we evaluate the SVG encoder against the DINO encoder on downstream tasks where DINO is known to excel. For each task, we adopt a simple strategy: a lightweight MLP- or linear-layer decoder is appended to the encoder to map features into predictions, and only the decoder is trained (details in Appendix B). As reported in Table 4, the SVG feature maintains the strong generalization ability of DINO, achieving comparable or even slightly superior results on ImageNet-1K (Deng et al., 2009; Russakovsky et al., 2015) classification, ADE20K (Zhou et al., 2019) semantic segmentation, and NYUv2 (Nathan Silberman and Fergus, 2012) depth estimation. Combined with its previously demonstrated strength in generative tasks, this dual advantage establishes the feature space produced by  SVG Encoder as a unified representation space for diverse vision tasks.
@@ 5 Conclusion
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62321005, Grant 62336004, Grant 62125603, and in part by the Beijing Natural Science Foundation under Grant No. L247009.
@@ Appendix A More implementation details
Classifier-free guidance. In the main text, we report results both with and without classifier-free guidance. To reduce the uncertainty in the initial velocity prediction, we adopt zero-init (Fan et al., 2025), which skips the first step. We report the FID50K with cfg 1.5 in our paper.
@@ Appendix B Finetuning Details on Downstream Tasks
We evaluate the SVG encoder against the DINOv3 encoder on three representative downstream tasks: ImageNet-1K (Deng et al., 2009; Russakovsky et al., 2015) classification, ADE20K (Zhou et al., 2019) semantic segmentation, and NYUv2 (Nathan Silberman and Fergus, 2012) depth estimation. For all tasks, the encoder remains frozen, and only lightweight decoders are trained.
Image classification. This task requires assigning each image to a single class. We report Top-1 and Top-5 accuracies. A linear classifier is placed on top of the encoder output to map features to class scores. Input images are randomly resized and cropped to $256\times 256$. The classifier is trained for 30 epochs with the AdamW optimizer (Loshchilov and Hutter, 2017), using a global batch size of 3072, an initial learning rate of 5e45e-4, and weight decay of 1e21e-2.
Semantic segmentation. The goal of semantic segmentation is to produce dense per-pixel predictions. We use mean Intersection-over-Union (mIoU) and mean Accuracy (mAcc) as evaluation metrics. The decoder adopts an FPNHead implementation in mmsegmentation, applied to the single-scale encoder feature. The training strategy generally follows (Zhao et al., 2023). Images are randomly resized and cropped to $512\times 512$ before being fed to the network. Optimization is performed with AdamW (Loshchilov and Hutter, 2017) at a learning rate of 8e58e-5, weight decay of 1e31e-3, and 1,500 warm-up steps. A polynomial scheduler with power 0.9 and a minimum learning rate of 1e61e-6 is used. Training runs for 8,000 iterations. During inference, we adopt sliding-window evaluation with $512\times 512$ crops and a stride of $341\times 341$.
Depth estimation. Depth estimation aims to regress pixel-wise depth values for input images. We report Absolute Relative Error (A.Rel) and RMSE as evaluation metrics. During training, images are randomly cropped to $480\times 480$. The model is optimized for 25 epochs using the AdamW (Loshchilov and Hutter, 2017) optimizer with a batch size of 24 and a learning rate of 5e45e-4. The decoder head and other hyperparameters follow the setup in (Xie et al., 2022a). At test time, we use both horizontal flipping and sliding-window inference.